OpenStudy (anonymous):

A university took a phone sample from 3421 students about tuition and found that 2004 of those students believed that tuition was too high. What is a good prediction of how many students think tuition is too high if there are about 20,000 students at the university? A. 11,616 students B. 11,516 students C. 11,716 students D. 11,816 students

OpenStudy (anonymous):

that is proportional, so... 3421 -> 2004 20000 -> x just solve the proportion and find x

OpenStudy (anonymous):

ok @M4thM1nd

OpenStudy (anonymous):

can u explain im blonde if u know what i mean @M4thM1nd

OpenStudy (anonymous):

Of course i can, and the color of your hair does not have any relationship with your intelligence :) btw i'm also blonde

OpenStudy (anonymous):

lol ok well im dumb then

OpenStudy (anonymous):

x = 17105000/501

OpenStudy (anonymous):

If something is proportional then there exists a constant C such that, in this case: C = 2004/3421, ok? So... this C values must also be equal here: C = x/20,000

OpenStudy (anonymous):

??

OpenStudy (anonymous):

is it right??

OpenStudy (anonymous):

No, that is not the correct answer.

OpenStudy (anonymous):

aww

OpenStudy (anonymous):

C = 2004/3421, and C = x/20,000, so... 2004/3421 = x/20000 x = 2004*20000/3421 = ...?

OpenStudy (anonymous):

UGH im soo stupid ok so 2004*20000=40080000

OpenStudy (anonymous):

Calculating Multipliers : 1.2 Calculate multipliers for the two fractions Denote the Least Common Multiple by L.C.M Denote the Left Multiplier by Left_M Denote the Right Multiplier by Right_M Denote the Left Deniminator by L_Deno Denote the Right Multiplier by R_Deno Left_M = L.C.M / L_Deno = 5000 Right_M = L.C.M / R_Deno = 501 Making Equivalent Fractions : 1.3 Rewrite the two fractions into equivalent fractions Two fractions are called equivalent if they have the same numeric value. For example : 1/2 and 2/4 are equivalent, y/(y+1)2 and (y2+y)/(y+1)3 are equivalent as well. To calculate equivalent fraction , multiply the Numerator of each fraction, by its respective Multiplier. L. Mult. • L. Num. 3421 • 5000 —————————————————— = ——————————— L.C.M 10020000 R. Mult. • R. Num. x • 501 —————————————————— = ———————— L.C.M 10020000 Adding fractions that have a common denominator : 1.4 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: 3421 • 5000 - (x • 501) 17105000 - 501x ——————————————————————— = ——————————————— 10020000 10020000 Equation at the end of step 1 : 17105000 - 501x ——————————————— = 0 10020000 Step 2 : 17105000-501x Solve ————————————— = 0 10020000 When a fraction equals zero : 2.1 When a fraction equals zero ... Where a fraction equals zero, its nominator, the part which is above the fraction line, must equal zero. Now,to get rid of the denominator, Tiger multiplys both sides of the equation by the denominator. Here's how: 17105000-501x ————————————— • 10020000 = 0 • 10020000 10020000 Now, on the left hand side, the 10020000 cancels out the denominator, while, on the right hand side, zero times anything is still zero. The equation now takes the shape : 17105000-501x = 0 Solving a Single Variable Equation : 2.2 Solve : -501x+17105000 = 0 Subtract 17105000 from both sides of the equation : -501x = -17105000 Multiply both sides of the equation by (-1) : 501x = 17105000 Divide both sides of the equation by 501: x = 17105000/501 One solution was found : x = 17105000/501

OpenStudy (anonymous):

didnt i do it right

OpenStudy (anonymous):

LOL there are a lot of interrogation marks in your message. i don't understand what you did

OpenStudy (anonymous):

and yes, 2004*20000=40080000

OpenStudy (anonymous):

it was a glitch or something lol

OpenStudy (anonymous):

now x = 40080000/3421

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Conqueror

OpenStudy (anonymous):

@King.Void.

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

i'm back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you understand the question now?

OpenStudy (anonymous):

no not realy

OpenStudy (anonymous):

i think its C though

OpenStudy (anonymous):

Yes it's C

OpenStudy (anonymous):

realy wow

OpenStudy (anonymous):

From 3421 students, 2004 believe tuition is high, so now, from 20,000 students, x believe tuition is high, so: 2004/3421 = x/20000 x = 11716

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

A candidate took a phone poll of 110 people. Of the 110 people polled, 87 said they would vote for the other person. There are 9570 people in the district. About how many people are going to vote for the other candidate? A. 7400 people B. 7569 people C. 7069 people D. 7900 people

OpenStudy (anonymous):

@M4thM1nd help me;]

OpenStudy (anonymous):

Same way, that is a proportion, C = 87/110, but C = x/9570 so... 87/110 = x/9570

OpenStudy (anonymous):

i think A ??

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

No... 87/110 = x/9570 x = 87*9570/110 = 832590/110

OpenStudy (anonymous):

ohhh so d

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@M4thM1nd i sent it 2 u